# Optimistic Verifiable Training by Controlling Hardware Nondeterminism

### v2

@@ Abstract
The increasing compute demands of AI systems has led to the emergence of services that train models on behalf of clients lacking necessary resources. However, ensuring correctness of training and guarding against potential training-time attacks, such as data poisoning and trapdoors, poses challenges. Existing works on verifiable training largely fall into two classes: proof-based systems, which can be difficult to scale, and “optimistic” methods that consider a trusted third-party auditor who replicates the training process. A key challenge with the latter is that hardware nondeterminism between GPU types during training prevents an auditor from replicating the training process exactly, and such schemes are therefore non-robust. We propose a method that combines training in a higher precision than the target model, rounding after intermediate computation steps, and storing rounding decisions based on an adaptive thresholding procedure, to successfully control for nondeterminism. Across three different NVIDIA GPUs (A40, Titan XP, RTX 2080 Ti), we achieve exact training replication at FP32 precision for both full-training and fine-tuning of ResNet-50 (23M) and GPT-2 (117M) models. Our verifiable training scheme significantly decreases the storage and time costs compared to proof-based systems.
Machine Learning, ICML
@@ 1 Introduction
Figure 1: Overview of our verifiable training scheme, based on Teutsch & Reitwießner (2019). After an $\mathrm{auditor}$ challenges a $\mathrm{trainer}$ on behalf of a client, they train the model themselves, storing weights in a Merkle tree, and enter a binary search procedure to identify the exact steps of the dispute. We show how to account for hardware nondeterminism so that the $\mathrm{auditor}$ and $\mathrm{trainer}$ can use different GPUs, expanding the pool of potential auditors to any party capable of performing the training task.
@@
Consider an education start-up that wishes to finetune the Llama-70b language model (70B parameters) on their own curated dataset to support student learning. This task requires significant resources, and the company might even lack the necessary expertise. Instead, they might choose to pay a $\mathrm{trainer}$ with vast computing resources to perform the training task (Figure 1). However, what if the $\mathrm{trainer}$ adds data points that spread misinformation, or introduces a backdoor that advances a political agenda for specific prompts? If the client starts to notice suspicious model behavior, is there any action they can take? We study this problem of verifiable training, or ensuring that the training of an ML model was performed correctly.
@@
An alternative “optimistic” approach is to consider a trusted third-pary $\mathrm{auditor}$, such as a non-profit organization, that has sufficient computing resources to perform the training task, even if not at the bandwidth of a service provider (Figure 1). When a $\mathrm{client}$ suspects foul play, they can ask the $\mathrm{auditor}$ to challenge the $\mathrm{trainer}$ by training the model using the $\mathrm{auditor}$’s own compute, and demonstrate that the $\mathrm{trainer}$ did not train correctly. This protocol can be efficiently carried out using techniques from the literature on verifiable computing, such as the “verification game” method of Teutsch & Reitwießner (2019). This method uses an interactive binary-search procedure to identify the exact intermediate computation step (e.g. training epoch) where the two parties diverged.
@@
In this paper we address this limitation by asking: can the $\mathrm{trainer}$ provide additional information to the $\mathrm{auditor}$ that removes the effects of hardware nondeterminism? Our starting point is the observation that hardware nondeterminism occurs due to the accumulation of errors from floating point operations. For example, a matrix-vector multiply often results in different floating point values on different GPUs. A natural approach is to perform training using a higher precision (e.g. FP32) than the target precision of the model weights (e.g. FP16), and periodically round back to the target precision. The hope is that all floating point errors will be confined to the higher precision bits, so that rounding will “clean” the computed values and ensure that the trainer and auditor stay in sync. However, this does not work. The reason is that the computed values occasionally straddle the rounding boundary, causing the $\mathrm{trainer}$ to round up while the $\mathrm{auditor}$ rounds down, quickly causing them to diverge. Alternatively, we propose a solution where the $\mathrm{trainer}$ records the rounding direction for every intermediate computation so that $\mathrm{auditor}$ can stay in sync with the $\mathrm{trainer}$. However, as this requires the $\mathrm{trainer}$ to record and output a large number of bits, we also show how to reduce the amount of data needed to eliminate errors due to hardware nondeterminism, as needed for correct auditing.
@@
We use this strategy to adapt the verification game described by Teutsch & Reitwießner (2019) for verifiable training. The game’s efficiency lies in our ability to store hashes of model checkpoints in a Merkle tree (Merkle, 1988). To determine if training was performed according to the specification, the $\mathrm{auditor}$ needs to reconstruct the Merkle tree and compare the resulting Merkle root hash with the Merkle root hash provided by the $\mathrm{trainer}$’s – if they do not match, the two parties can enter an interactive binary search procedure to identify the exact training step of the dispute. The purpose of the binary search game is to hold both parties accountable: an $\mathrm{auditor}$ should not be able to simply claim that a model was improperly trained, but convince a third-party (e.g., the public, or a $\mathrm{judge}$) by showing at what point during training the $\mathrm{trainer}$ misbehaved. We show this verifiable training scheme can scale to tasks such as full training of ResNet-50 (23M parameters) and finetuning of GPT-2 (117M parameters), significantly outperforming existing methods with respect to both time and storage cost, while eliminating statistical error due to non-determinism. For example, the proposal in prior work Jia et al. (2021) would require $\mathbf{>140\times}$ more storage cost than our method by saving and comparing model weights at every step in order to achieve the lowest possible (yet still non-zero) statistical error.
@@
Concretely, our contributions include: (1) A method for two parties, training the same model on different GPU types, to achieve identical results based on sharing rounding decisions; (2) A verifiable training scheme based on the verification game from Teutsch & Reitwießner (2019), which stores model weights in a Merkle tree for efficient comparison between a $\mathrm{trainer}$ and $\mathrm{auditor}$; (3) Experiments showing the ability of our approach to scale to large models such as ResNet-50 and GPT-2 between three different NVIDIA GPU architectures (A40, Titan XP, RTX 2080 Ti); (4) Methods to reduce the storage cost of our approach via efficient encoding of rounding logs and an adaptive threshold mechanism to reduce the amount of rounding decisions logged; and (5) Comparisons with existing methods, including proof-based systems, that highlight the improved storage and time efficiency of our method.  [^1] Our method is implemented entirely within the pytorch framework, and is available at https://github.com/meghabyte/verifiable-training.
@@ 2 Related Works
Attacks on Model Training: Without any verifiable training scheme, significant trust is placed in the $\mathrm{trainer}$. This leaves a client vulnerable to a variety of different attacks, such as “poisoning” of data samples to cause undesirable model behaviors (e.g., generating unsafe, text (Carlini et al., 2023; Koh et al., 2021; Wan et al., 2023)) and planting of hard to detect & remove backdoors triggered by certain inputs (Hubinger et al., 2024; Goldwasser et al., 2022).
@@
Verifiable ML: Training and running inference on ML models in trusted environments has been an exciting direction explored by many researchers. One line of work consists of proof-based systems, where a proof of correctness (for a desired specification) is provided using cryptographic techniques such as succinct non-interactive arguments (SNARKs) (Micali, 1994; Bitansky et al., 2012; Lee et al., 2020; Liu et al., 2021; Garg et al., 2023; Kang et al., 2022). The majority of these works focus on the easier problem of verifiable inference, or proving that a prediction indeed came from a given model. However, such methods require approximations for non-linear functions such as softmax and ReLU activations, and require proof times on the order of minutes for a single data input (Liu et al., 2021). Subsequently, even the most recent proof-based systems for verifiable training suffer extreme latency, such as 22 minutes for training VGG-11 on one batch of 16 data inputs (Abbaszadeh et al., 2024), and efficient protocols have only been developed for simple models such as linear and logistic regression, for which it is less likely for a client to delegate out training in the first place (Garg et al., 2023; Ames et al., 2022).
An alternative direction is to consider hardware-based solutions, such as training models in a trusted execution environment (TEE). Such approaches would only require a 3rd-party $\mathrm{auditor}$ to review the source code and training data that is used by the $\mathrm{trainer}$ and attested for by the TEE, and not perform expensive model training themselves as in our approach. However, TEEs such as NVIDIA’s H100 “Confidential GPU” incur a performance penalty for model training due to the cost of running inside a TEE (Dhanuskodi et al., 2023). Moreover, clients run the risk of losing all security guarantees if an attacker can extract the attestation key from even one GPU, as has been successfully demonstrated in several works (Nilsson et al., 2020; Bulck et al., 2018).
@@
Most similar to our approach are proof-of-learning protocols, which also consider a trusted 3rd party that compares checkpointing during the course of training with the original training sequence (Jia et al., 2021). However, such methods not only incur high storage cost by requiring model weights to be stored frequently, but are also subject to errors from training nondeterminism, and cannot provide a robust guarantee. Several works have shown that proof-of-learning protocols can be spoofed and fail to verify correctness in several important contexts, including training tasks such as machine unlearning, where the $\mathrm{trainer}$ needs to produce a new model that “forgets” specific training data due to privacy legislation (Fang et al., 2023; Kong et al., 2023; Thudi et al., 2022). Although Choi et al. (2023) recently proposed a verification procedure that is immune to several known proof-of-learning attacks, their method is not only limited to supervised learning algorithms, but also based on an assumption that models temporarily overfit data during training, which may not hold true for all training tasks. Overall, these works argue that the main limitation for provably robust protocols is error due to training nondeterminism. This provides further motivation for our work.
@@
GPU Nondeterminism: Prior work has investigated software patches for deterministic training, for instance by enforcing floating point accumulation ordering, at a significant cost to efficiency (Jooybar et al., 2013; Defour & Collange, 2015; Chou et al., 2020; TensorFlow, 2021; Zhuang et al., 2021). However, while these options address the issue of deterministic computation on a single GPU architecture — achieving deterministic results across multiple GPU architectures remains challenging (Crane, 2018a; NVIDIA, 2022). Our work seeks to control hardware nondeterminism across architecture types in order to design an efficient and reliable verifiable training scheme. The value of a deterministic training scheme extends beyond verifiable training — researchers have shown that training nondeterminism can have several negative consequences, including concerns of fairness due to causing inconsistent performance on minority subgroups of the data, the reproducibility of published results, and downstream effects on complex ML pipelines (Zhuang et al., 2021; Crane, 2018b; Srivastava et al., 2020).
@@ 3 Set-Up: The Verification Game
Figure 2: Even after controlling for the same software version, random seed, and use of nondeterministic algorithms via library flags, training nondeterminism persists between GPU types, posing a challenge for verifying training correctness.
@@
Our method for verifiable training is based on the interactive verification game proposed by Teutsch & Reitwießner (2019) in the context of blockchains. The basic idea of the verification game is to resolve a dispute between a challenger, in our case the $\mathrm{auditor}$, and a solver, in our case the $\mathrm{trainer}$, for an expensive computation (e.g. model training). In order for the $\mathrm{auditor}$ to take any meaningful action (e.g. pursue legal action), they need to prove the exact source of the dispute (e.g. training time-step where an attack occurred). If we can save model weights at different time steps into a compact data structure such as a Merkle tree, then identifying the source of disagreement can be done efficiently using binary search (Merkle, 1988). More precisely, the verification game consists of the following parties:
@@
  • $\mathrm{trainer}$, who has putatively trained a model according to a set of specifications. In our example, this is a service provider (e.g. a large company) with sufficient compute power to train a model on behalf of a client.
  • $\mathrm{client}$, who receives a model from the $\mathrm{trainer}$ and is concerned about its correctness. They are unable to perform the training task themselves, and are willing to share their data with both $\mathrm{trainer}$ and $\mathrm{auditor}$.
@@
  • $\mathrm{auditor}$, who officially challenges the $\mathrm{trainer}$ on behalf of a client. This is a trusted 3rd-party that also has sufficient resources to perform the same training task as the $\mathrm{trainer}$, but does not provide such training as a service. If needed, the client can choose several auditors to audit the model that was generated by the $\mathrm{trainer}$. While an $\mathrm{auditor}$ could be a trusted regulatory organization, it could also be some other provider of training services.
@@
  • $\mathrm{judge}$: In some cases, it may be desirable to pursue legal action and have a $\mathrm{judge}$ arbitrate. The $\mathrm{judge}$ can only perform minimal computations (e.g. training for one epoch), but can examine the $\mathrm{auditor}$’s claims and enforce a penalty against either the $\mathrm{trainer}$, for training incorrectly, or the $\mathrm{auditor}$, for raising a false alarm.
@@
When the $\mathrm{trainer}$ is approached by an $\mathrm{auditor}$, they would need to share training parameters, model architecture, and randomness, as shown in Figure 1. The $\mathrm{auditor}$ would then replicate the training process, storing model weights in a Merkle tree at the same checkpoints as the trainer. Recall that a Merkle tree is a tree where every leaf node is a hash of the model weights and every non-leaf node is a hash of its children nodes. The main loop of the verification game starts when both parties have the root hash of their respective Merkle trees. If training was performed correctly, then the $\mathrm{trainer}$’s root should match the $\mathrm{auditor}$’s. Otherwise, a binary search procedure is performed, where the $\mathrm{auditor}$ iteratively descends the Merkle tree until it identifies two consecutive leaf nodes, $i$ and $i+1$, where the hash at $i$ matches that of the $\mathrm{trainer}$, but the hash at leaf $i+1$ does not. This identifies the point in the computation that caused the dispute.
@@
This interactive verification game requires the cooperation of the $\mathrm{trainer}$. If the $\mathrm{trainer}$ refuses to share the value at a certain node of their Merkle tree within a given time frame, they can be considered to have failed the audit. Additionally, we emphasize that the $\mathrm{trainer}$ and $\mathrm{auditor}$ use a Merkle tree to store model weights, requiring far less storage than prior work, if correct training produces identical weights (and therefore identical hash values).
The problem is that training nondeterminism between GPUs leads to weight divergence between the auditor and trainer, and causes the verification game outlined above to always fail. This why we seek to prevent divergence in training.

### v3

@@ Abstract
The increasing compute demands of AI systems have led to the emergence of services that train models on behalf of clients lacking necessary resources. However, ensuring correctness of training and guarding against potential training-time attacks, such as data poisoning and backdoors, poses challenges. Existing works on verifiable training largely fall into two classes: proof-based systems, which are difficult to scale, and “optimistic” methods that consider a third-party auditor who can replicate the training process and contest the trainer. A key challenge with the latter is that nondeterminism between GPU types during training prevents exact replication of the training process, resulting in schemes that are non-robust. We propose a method that combines training in a higher precision than the target, rounding after intermediate computations, and sharing rounding decisions based on an adaptive thresholding procedure, to successfully control for nondeterminism. Across three different NVIDIA GPUs (A40, Titan XP, RTX 2080 Ti), we achieve exact training replication at FP32 precision for both full-training and fine-tuning of ResNet-50 (23M) and GPT-2 (117M) models. Our verifiable training scheme significantly decreases the storage and time costs compared to proof-based systems, and is publicly released at https://github.com/meghabyte/verifiable-training.
@@ 1 Introduction
Figure 1: Overview of our scheme. After an $\mathrm{auditor}$ challenges a $\mathrm{trainer}$, they train the model, storing weights in a Merkle tree, and enter a binary search procedure to identify the exact steps of the dispute. We show how to control GPU nondeterminism between $\mathrm{auditor}$ and $\mathrm{trainer}$, expanding the set of potential auditors.
@@
Consider an education start-up that wishes to finetune the Llama-70b language model (70B parameters) on their own curated dataset to support student learning. This task requires significant resources, and the company might even lack the necessary expertise. Instead, they might choose to pay a $\mathrm{trainer}$ with vast computing resources to perform the training task (Figure 1). However, what if the $\mathrm{trainer}$ adds data points that spread misinformation, introduces a backdoor that advances a political agenda for specific prompts, or tries to save work by under-training the model? If the client starts to notice suspicious model behavior, is there any action they can take? We study this problem of verifiable training, or ensuring that the training of an ML model was performed correctly.
@@
An alternative “optimistic” approach is to consider a third-pary $\mathrm{auditor}$ (Figure 1). This could be a trusted 3rd party, such as a non-profit organization that may not have sufficient computing resources to provide training as a service beyond auditing, or a different provider that the client approaches and wishes to compare with the original model trainer. When a $\mathrm{client}$ suspects foul play, they can ask the $\mathrm{auditor}$ to challenge the $\mathrm{trainer}$ by training the model using the $\mathrm{auditor}$’s own compute, and demonstrate that the $\mathrm{trainer}$ did not train correctly. Based on the provided evidence required from the $\mathrm{auditor}$ (i.e. the precise timesteps model training diverged, as shown in Figure 1), the $\mathrm{client}$ can then choose to refuse the $\mathrm{trainer}$’s model, pursue legal action against the $\mathrm{trainer}$, or even dispute a potentially corrupt $\mathrm{auditor}$ if the client deems such evidence as incorrect, or another $\mathrm{auditor}$ disagrees. This protocol can be efficiently carried out using techniques from the literature on verifiable computing, such as the “verification game” method of Teutsch and Reitwießner (2019), which uses an interactive binary-search procedure to identify the exact intermediate computation step (e.g. training epoch) where the two parties diverged. Applying verifiable computation techniques to model training is particularly important given the increase in decentralized machine learning services like Gensyn, which seek to make ML compute more accessible by creating a network of many untrusted GPUs.
@@
Our work addresses this limitation by asking: can the $\mathrm{trainer}$ provide additional information to the $\mathrm{auditor}$ that removes the effects of hardware nondeterminism? Our starting point is the observation that hardware nondeterminism occurs due to the accumulation of errors from floating point operations. For example, a matrix-vector multiply often results in different floating point values on different GPUs, since GPUs often accumulate in different orders. To address this issue, a natural approach is to perform training using a higher precision (e.g. FP32) than the target precision of the model weights (e.g. FP16), and periodically round back to the target precision. The hope is that all floating point errors will be confined to the higher precision bits, so that the rounded values are deterministic. However, this fails because computed values can occasionally straddle the “rounding boundary”: i.e., the $\mathrm{trainer}$ can round up while the $\mathrm{auditor}$ rounds down, quickly causing them to diverge. Instead, we propose a solution where the $\mathrm{trainer}$ records the rounding direction for certain intermediate computation so that $\mathrm{auditor}$ can stay in sync with the $\mathrm{trainer}$. As this requires the $\mathrm{trainer}$ to record a large number of bits, we also show how to reduce the amount of data needed to eliminate errors.
@@
We use this strategy to adapt the verification game described by Teutsch and Reitwießner (2019) for verifiable training. The game’s efficiency lies in our ability to store hashes of model checkpoints in a Merkle tree (Merkle, 1988). To determine if training was performed according to the specification, the $\mathrm{auditor}$ needs to reconstruct the Merkle tree and compare the resulting Merkle root hash with the Merkle root hash provided by the $\mathrm{trainer}$’s – if they do not match, the two parties can enter an interactive binary search procedure to identify the exact training step of the dispute. The purpose of the binary search game is to hold both parties accountable: an $\mathrm{auditor}$ should not be able to simply claim that a model was improperly trained, but convince a third-party (e.g., the public, or a $\mathrm{judge}$) by showing at what point during training the $\mathrm{trainer}$ misbehaved. We show our verifiable training scheme can scale to tasks such as full training of ResNet-50 (23M parameters) and finetuning of GPT-2 (117M parameters), significantly outperforming existing methods with respect to both time and storage cost, while eliminating statistical error due to non-determinism. For example, the proposal in prior work Jia et al. (2021) would require $\mathbf{>140\times}$ more storage cost than our method by comparing model weights at every step in order to achieve low (yet still non-zero) statistical error.
@@
Concretely, our contributions include: (1) A method for two parties, training the same model on different GPU types, to achieve identical training results by logging and sharing rounding decisions; (2) A verifiable training scheme based on the verification game from Teutsch and Reitwießner (2019), which stores model weights in a Merkle tree for efficient comparison between a $\mathrm{trainer}$ and $\mathrm{auditor}$; (3) Experiments showing the ability of our approach to scale to large models such as ResNet-50 and GPT-2 between three different NVIDIA GPU architectures (A40, Titan XP, RTX 2080 Ti); (4) Methods to reduce the storage cost of our approach via efficient encoding of rounding logs and an adaptive threshold mechanism to reduce the amount of rounding decisions logged; and (5) Comparisons with existing methods, including proof-based systems, that highlight the improved storage and time efficiency of our method.  [^1] Our method is implemented entirely within the pytorch framework (compatible with version 2.3.1), and is available at https://github.com/meghabyte/verifiable-training.
@@ 2 Related Works
Without any verifiable training scheme in place, significant trust is placed in the $\mathrm{trainer}$, leaving a client vulnerable to many different attacks, such as the “poisoning” of data samples to cause undesirable behavior (e.g., generating unsafe text (Carlini et al., 2023; Koh et al., 2021; Wan et al., 2023)) and planting backdoors triggered by certain inputs (Goldwasser et al., 2022). Therefore, training ML models in trusted environments has been an exciting direction explored by many researchers. One line of work consists of proof-based systems, where a proof of correctness (for a desired specification) is provided using cryptographic techniques such as succinct non-interactive arguments (SNARKs) (Micali, 1994; Bitansky et al., 2012; Lee et al., 2020; Liu et al., 2021; Garg et al., 2023; Kang et al., 2022). However, even the most recent proof-based systems for verifiable training suffer extreme latencies, such as 22 minutes for training VGG-11 on one batch of 16 data inputs (Abbaszadeh et al., 2024), and have therefore primarily been developed for simpler models (e.g., logistic regression) that are less likely to be delegated to others in the first place (Garg et al., 2023; Ames et al., 2022). Meanwhile, an alternative solution of training models in a trusted execution environment (TEE), such as NVIDIA’s H100, incurs a performance penalty due to the cost of running inside a TEE (Dhanuskodi et al., 2023). Furthermore, clients lose all security guarantees if an attacker can extract the attestation key from even one GPU  (Nilsson et al., 2020; Bulck et al., 2018).
@@
Our approach is most similar to proof-of-learning protocols, which consider a trusted 3rd party that compares checkpointing during the course of training with the original training sequence (Jia et al., 2021). However, such methods not only incur high storage cost by requiring model weights to be stored frequently, but are non-robust due to errors from training nondeterminism. Several works have shown that proof-of-learning protocols can be spoofed and fail to verify correctness in several important contexts (Fang et al., 2023; Kong et al., 2023; Thudi et al., 2022). Although Choi et al. (2023) recently proposed a verification procedure that is immune to several known proof-of-learning attacks, their method is not only limited to supervised learning algorithms, but also based on an assumption that models temporarily overfit data during training, which may not always hold true.
@@
GPU Nondeterminism: Prior work has investigated software patches for deterministic training, for instance by enforcing FP accumulation ordering, at a significant cost to efficiency Jooybar et al. (2013); Defour and Collange (2015); Chou et al. (2020); TensorFlow (2021); Zhuang et al. (2021). While these options address deterministic computation on a single GPU architecture, achieving deterministic results across multiple GPU architectures remains challenging Crane (2018a); NVIDIA (2022). We control hardware nondeterminism across GPUs in order to design an efficient and reliable verifiable training scheme. However, our method’s impact extends beyond verifiable training, as training nondeterminism can have several negative consequences including bias, reproducibility, and downstream effects on ML pipelines (Zhuang et al., 2021; Crane, 2018b; Srivastava et al., 2020).
@@ 3 Set-Up: The Verification Game
Figure 2: Even after ensuring the same software version, random seed, and use of deterministic algorithms via library flags, training nondeterminism persists between three GPU types.
@@
Our method for verifiable training is based on the interactive verification game proposed by Teutsch and Reitwießner (2019) in the context of blockchains. The core idea is to resolve a dispute between a challenger, in our case the $\mathrm{auditor}$, and a solver, in our case the $\mathrm{trainer}$, for an expensive computation (e.g., model training). In order for the $\mathrm{auditor}$ to take any meaningful action (e.g., pursue legal action), they need to prove the exact source of the dispute (e.g., training time-step where an attack occurred). If we can save model weights at different time steps into a compact data structure such as a Merkle tree, then identifying the source of disagreement can be done efficiently using binary search (Merkle, 1988). More precisely, the verification game consists of the following parties:
  1. 1.
@@
    $\mathrm{trainer}$, who has putatively trained a model according to a $\mathrm{client}$’s specifications. In our example, this is a service provider with sufficient compute power to train a model.
  2. 2.
    $\mathrm{client}$, who receives a model from the $\mathrm{trainer}$ and approaches an $\mathrm{auditor}$.
  3. 3.
@@
    $\mathrm{auditor}$, who officially challenges the $\mathrm{trainer}$ on behalf of a client. This is a trusted 3rd-party that has sufficient resources but does not necessarily provide training as a service. The $\mathrm{client}$ can choose several auditors to audit the $\mathrm{trainer}$’s model.
  4. 4.
@@
    $\mathrm{judge}$: Sometimes a $\mathrm{judge}$ may need to arbitrate a legal claim. The $\mathrm{judge}$ can only perform minimal computations (e.g., one training epoch), but can examine the $\mathrm{auditor}$’s claims and enforce a penalty against either the $\mathrm{trainer}$, for incorrect training, or the $\mathrm{auditor}$, for a false alarm.
@@
When the $\mathrm{trainer}$ is approached by an $\mathrm{auditor}$, they would need to share training parameters, model architecture, and randomness, as shown in Figure 1. The $\mathrm{auditor}$ would then replicate the training process, storing model weights in a Merkle tree at the same checkpointing interval as the trainer (every leaf node in a Merkle tree is a hash of the data and every non-leaf node is a hash of its children). The main loop of the verification game starts when both parties have the root of their respective Merkle trees. If training was performed correctly, then the $\mathrm{trainer}$’s root should match the $\mathrm{auditor}$’s. Otherwise, a binary search procedure is performed, where the $\mathrm{auditor}$ iteratively descends the Merkle tree until it identifies two consecutive leaf nodes, $i$ and $i+1$, where the hash at $i$ matches that of the $\mathrm{trainer}$, but the hash at leaf $i+1$ does not. This identifies the point in the computation of the dispute.
@@
This interactive verification game requires the cooperation of the $\mathrm{trainer}$. If the $\mathrm{trainer}$ refuses to share the value at a certain node of their Merkle tree within a given time frame, they can be considered to have failed the audit. Additionally, the $\mathrm{trainer}$ and $\mathrm{auditor}$ use a Merkle tree to store model weights, requiring far less storage than prior work, if correct training produces identical weights (and identical hash values).The problem is that training nondeterminism leads to weight divergence, and causes this verification game to always fail. This why we seek to prevent divergence in training.